OpenStudy (anonymous):

what is the standard notation of 5.26 * 10^-4 and 3.18 * 10^3

OpenStudy (campbell_st):

ok.... so you have your number in scientific notation so its simply \[\frac{5.26}{10000} = ?\] because 10^-4 is the same as dividing by 10000 do the calculation to get an answer or ? the other question is 3.18 x 1000 since 10^3 = 1000

OpenStudy (anonymous):

\[5.26 \times 10^{-4} = \frac{ 5.26 }{ 10000 } = 0.000526\]

OpenStudy (campbell_st):

thats correct.... its really just a case of moving the decimal point left or right... hope it helped